# Exploring the Interior Structure and Mode of Tidal Heating in Enceladus

### v3

@@ 6 Road-map to Constrain the interior properties
Despite numerous studies based on Cassini observations, numerous interior structure properties of Enceladus are poorly constrained. This uncertainty is due in part to the limitation of the available observations and in part to differences in the interpretation of the data for the interior structure and the underlying modeling assumptions (e.g., Howett et al., 2011; Spencer et al., 2013; Thomas et al., 2016; Park et al., 2024; Schenk & McKinnon, 2024). Here, we explore the ranges of possible interior parameters that are consistent with the currently available observations. Based on these constraints, we predict ranges of the currently unavailable geodetic observables, and investigate the possibility of discriminating between the two main tidal heating scenarios from future measurements. We use the estimates of the observables as a basis for exploring the interior from future measurements.
6.1 Constraining the Interior and Discriminating Between Dissipation Scenarios
@@
Currently available observations that can be used to understand Enceladus’s interior structure and heat production include shape (e.g., Park et al., 2024; Schenk & McKinnon, 2024), gravity up to degree-3 (Iess et al., 2014; Park et al., 2024), physical libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024), and heat flux (Howett et al., 2011; Spencer et al., 2013) as summarized in Table 1. We highlight that in this paper, in the absence of more evidence, we assume that the total heat loss is equal to the tidally generated heat computed from Equation (5); in other words, we assume that Enceladus is currently in a thermal steady state. Moreover, we also note that the rates of heat loss considered here, except by Howett et al. (2011), are not based on direct measurements, but are estimated from the shell thickness and assumptions on its thermal conductivity. The rate of heat loss is largely used to determine the tidal phase lag. To avoid any self-consistency issue, we ensure that the obtained a posteriori ice shell thicknesses are within the ranges used to determine the adopted rates of heat loss. We further discuss the rate of heat loss in Enceladus in Section 7.
@@
Figure 4: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and libration amplitude spanned by estimates in Thomas et al. (2016), Nadezhdina et al. (2016) and Park et al. (2024). Other observables in this plot are free to vary. In the 2D Probability Density Functions (PDFs), black (yellow) color indicates low (high) probability. The observed multiple peaks in the PDFs of $|k_{2}|$, $|h_{2}|$, $|l_{2}|$, and phase lags reflect the different regimes of dissipation consistent with the observations (see Figure 2), resulting in a nonlinear forward model. Phase lags of $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, and $\epsilon_{l_{2}}$ are in degrees, amplitude of the libration is in meters, and $J_{3}$ is divided by 106.
6.1.1 Plausible Interior Parameters and Observables
@@
Consistent with Section 3, the patterns of the posterior probability density functions (PDFs) of the amplitude and phase lag of $k_{2}$ largely differ from the PDFs of the amplitudes and phases of $h_{2}$ and $l_{2}$ (Figure 4). PDFs of $|k_{2}|$ and $\epsilon_{k_{2}}$ have several peaks associated with tidal activity dominated in the core versus in the shell. The observed peaks in this figure are associated with the ranges of the plausible dissipation mechanisms both in the shell and in the core that are depicted in Figure 1. A different distribution pattern and peaks associated with different regimes of tidal activity are observed for $h_{2}$ and $l_{2}$, and their phase lags. This different pattern arises from the fact that $|h_{2}|$ and $|l_{2}|$ are largely insensitive to tidal activity in the core. The posterior PDFs show that while the $|h_{2}|$ and $|l_{2}|$ and their phase lags (Figure 4) are strongly correlated with each other (dominated by the response of the shell), their correlation with the amplitude and phase lag of $k_{2}$ depends on where the tidal dissipation dominantly occurs.
@@
We also obtain the ranges of the interior model parameters and the correlation between them associated with the observables (Appendix E, Figure 18). Previously, we ruled out mean crustal thicknesses around 5–10 km for Enceladus, as they would result in amplitudes of libration that are too large (see Section 4). Ice shell thicknesses $\approx$ 1–5 km would be compatible with the estimates of tidal heating and libration only if the mean effective viscosity of the shell is unrealistically low. Therefore, we conclude that the mean thickness of Enceladus’s shell is larger than 10 km. For completeness, we include very low effective viscosities for the shell and the core, which imply deformation in the viscous-dominated regimes. As a result, the PDFs of the viscosities of the shell and core indicate multiple peaks representative of tidal dissipation dominated by viscous deformation as well as shear (anelastic) relaxation (See also Figure 1). In the remaining analysis below, we consider viscosities of the ice shell and the rocky core that are close to or greater than their nominal relaxation viscosities ($\eta_{\rm shell}~>~10^{12}$ Pa.s and $\eta_{\rm core}~>~10^{11}$ Pa.s), i.e., the ice shell and the rocky core deform in the anelasticity-dominated deformation regime and not in the viscous-dominated deformation regime.
@@
We focus on the analysis of discriminating between two sets of geodetic observables and interior parameters that are associated with each of the two scenarios (Figure 5). In this figure, and in the remainder of this paper, we distinguish between core-dominated and shell-dominated tidal dissipation scenarios by associating them with the cases where at least 90% of the total heat is produced in the shell or the core. We compute the tidal heating in the shell using Equation (14). The PDFs of $|k_{2}|$ and $\epsilon_{k_{2}}$ for the two tidal dissipation scenarios largely overlap, whereas the PDFs of $|h_{2}|$, $\epsilon_{h_{2}}$, $|l_{2}|$, and $\epsilon_{l_{2}}$ in the case of shell-dominated tidal dissipation differ significantly from those in the case of core-dominated tidal dissipation (Figure 5). In an inverse problem context, the degeneracy caused by the two different tidal dissipation scenarios reflected in the tidal Love numbers, as discussed earlier, leads to the non-uniqueness of the plausible interpretation of $k_{2}$ and $\epsilon_{k_{2}}$.
@@
By discriminating between the two tidal dissipation scenarios, we obtain a tight constraint on the viscosity of the shell in the cases where the tidal dissipation is dominated in the shell. Similarly, we constrain the viscosity of the core in the case where we assume that tidal dissipation occurs predominantly in the core (Figure 6). Based on the set of observables used in Figures 4 and 5, we find that the thickness of the ice shell, mostly constrained by the libration amplitude, is in the range of 15–40 km. Because we have conservatively chosen broad ranges for observables in this analysis, we have not obtained tight constraints on most of the other model parameters in this analysis. The PDFs and correlations between the interior structure parameters are presented in Appendix E.
@@
Figure 5: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux and the libration amplitude for tidal dissipation scenarios in the shell and in the core. The heat flux is assumed as $\rm 15~GW<\dot{E}<40~GW$, covering the estimated ranges in Hemingway et al. (2018); Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is conservatively assumed to be between $\rm 0.082^{\circ}<\gamma<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). In the 2D PDFs, the pink-gray and blue-yellow themes refer to dissipation in the shell and and in the core, respectively. All the variables are shown in logarithmic axes. Unit conventions are the same as in Figure 4.
@@
We compare the PDFs of $h_{2}$ and $l_{2}$ and their phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) associated with the two heat generation scenarios to compute the minimum required precision of measurements to discriminate between the two dissipation scenarios (Figure 5). We compute the minimum measurement precision of $|h_{2}|$ and $|l_{2}|$ by calculating the separation between the two intervals determined by the 2-$\sigma$ uncertainties associated with the PDFs. Therefore, if we measure $h_{2}$ with a precision equal to the separation between the 2-$\sigma$ confidence intervals (or 1-$\sigma$, not shown in the figure) of the two PDFs, we obtain a confidence $\sim$95% (or 68% if the 1-$\sigma$ intervals are used) to identify the layer that dominates the dissipation budget. We present more detailed comparison of $|h_{2}|$ and $|l_{2}|$ and their phase lags in Appendix E.
@@
Due to the overlap between the PDFs of $|h_{2}|$ and $|l_{2}|$ associated with the two tidal dissipation scenarios using this set of observations, distinguishing between the heating in the shell or in the core is challenging. This large overlap is partly due to the large uncertainties in the observables considered in this case, and taking into account only the currently available constraints. The analysis implies that measurement of the maximum vertical tidal deformations with a maximum uncertainty of 62 cm can distinguish between the dominant region of tidal activity within the 1-$\sigma$ confidence level. Measurement of radial deformations with a maximum uncertainty of 56 cm can indicates the “more likely” region of heat generation, but it cannot definitively discriminate between the two scenarios. Alternatively, measurement of the tangential displacements within an uncertainty of 28 cm is required for 1-$\sigma$ confidence. The phase lags associated with the radial deformations ($\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$) indicate a better separation between the PDFs associated with the two regimes of tidal activity compared to the deformations and $|h_{2}|$. We find that measuring the phase lag angles of the tidal deformations with an accuracy of $1$ is required for 1-$\sigma$ confidence, and 4 is required for 2-$\sigma$ confidence in determining the dominant region of tidal dissipation (see Appendix E for more details).
6.1.2 Measurement Requirements Based on Park et al. (2024)
@@
Next, we focus on a narrower range of observational constraints by exclusively using the ranges of $\gamma$, $J_{3}$, moment of inertia (inferred from $C_{20}$ and $C_{22}$), total rate of heat loss (inferred from conductive shell thickness), and the density based on the analysis by Park et al. (2024) (see Table 1). We use 3-$\sigma$ uncertainties on all the observables. We present the correlation between the observables obtained based on this analysis and incorporating both tidal dissipation scenarios in Figures 7. Applying the narrower observational constraints based on Park et al. (2024) shows that $|k_{2}|$ and $\epsilon_{k_{2}}$ indicate a normal distribution with a single peak that encompasses both scenarios of tidal dissipation, making them indistinguishable based only on measurement of $|k_{2}|$ and $\epsilon_{k_{2}}$. However, $|h_{2}|$, $|l_{2}|$, $\epsilon_{h_{2}}$, and $\epsilon_{l_{2}}$ show two peaks associated with the two tidal dissipation scenarios. By separating the two tidal dissipation regimes (See Appendix E,  Figures 18b and 18c), we also quantify the preferred values of the observables that present the highest likelihood for each of the two heat generation scenarios based on the analysis of Park et al. (2024) (Table 3). We use these synthetic observations to perform inversions using plausible uncertainties expected from future measurements. We note that satisfying all constraints provided in Park et al. (2024) is possible when using 3-$\sigma$ uncertainties on the observables. Specifically, we find that the mean value of $J_{3}$ can be matched with the libration amplitude when considering 3-$\sigma$ uncertainty in the measured $J_{3}$. Our results can be compared to those presented by Genova et al. (2024) and Ermakov et al. (2021), where only tidal heating in the shell is considered. Also, Genova et al. (2024) use only the gravity measurements without exploiting the libration measurement as a constraint, which is expected to somewhat alter the results.
@@
Table 4: Ranges of the model parameters from two of MCMC analysis satisfying the currently available measurements based on the analysis of observables in Park et al. (2024).
The tighter constraints obtained from Park et al. (2024) allow relaxation of the measurement requirement for $h_{2}$ and $l_{2}$ in order to discriminate between the two heating scenarios. Based on the observational constraints of Park et al. (2024), we show the amplitudes and phase lags of the Love numbers in Figure 8 panels (a)–(f). Using this narrower set of observational constraints, the PDFs of the tidal response associated with the two tidal dissipation scenarios are further separated compared to Section 6.1.1, where broader ranges of values for the observables were assumed. In this case, measuring the maximum vertical deformations with a precision of 87 cm is sufficient to differentiate between the two hypotheses of heat generation with a 1-$\sigma$ confidence. For a confidence level of 2-$\sigma$, measuring the maximum radial displacement of 1 cm is required. Alternatively, measuring the phase lag $\epsilon_{h_{2}}$ by 20 and 15 is sufficient to discriminate between the two tidal dissipation scenarios with 1-$\sigma$ and 2-$\sigma$ confidences, respectively. In addition, our analysis indicates that measurements of the maximum tangential displacements within 40 cm and 0.5 cm are required to differentiate between the two tidal dissipation scenarios providing 1-$\sigma$ and 2-$\sigma$ level confidences, respectively. Similarly to $\epsilon_{h_{2}}$, the measurement precisions required for $\epsilon_{l_{2}}$ for the confidence levels of 1-$\sigma$ and 2-$\sigma$ are 20 and 15, respectively.
In case future measurements provide constraints on $k_{2}$, the measurement requirement for $h_{2}$ and/or $l_{2}$ could be significantly relaxed. We present PDFs of displacement tidal Love numbers associated with the core- and shell-dominated tidal dissipation assuming constraints on the amplitude and phase lag of $k_{2}$ (Figure 8 panels (g)–(j)). If the real and imaginary parts of $k_{2}$ are measured within 0.002, the accuracies required to measure the maximum radial deformations to obtain 1-$\sigma$ and 2-$\sigma$ confidence levels relax to 134 cm and 82 cm, respectively. In that case, the upper bounds of the uncertainty required for the horizontal deformations are 61 cm and 36 cm, respectively. The accuracies required for the phase lags of $h_{2}$ or $l_{2}$ are equal to 21 and 16 for confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively.
Furthermore, in case more precise measurements of $k_{2}$ (both amplitude and phase) are obtained, i.e. within an uncertainty of 0.0002, the required measurement of radial or horizontal deformations will be further relaxed (not shown here). In that case, the requirement for the maximum radial displacement will be 150 cm and 116 cm for the 1-$\sigma$ and 2-$\sigma$ confidence levels, respectively. These values are 68 cm and 51 cm for the tangential component of the deformation, respectively. The precisions required for the phase lags of $h_{2}$ or $l_{2}$ are equal to 24 and $2$ to obtain confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively. We summarize the precision required for measuring the radial and horizontal displacements based on different scenarios discussed here in Table 5. Note that the measurements of the $h_{2}$ and $l_{2}$ as well as the $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ are correlated with each other and, therefore, measuring each of these four quantities is sufficient for discriminating between tidal dissipation scenarios. We perform our analysis based on both the derivation in this paper (Equations 16) and that in Beuthe (2019). We find that the difference in the results is less than 5%. This close approximation indicates the applicability of (Equations 16) in capturing the tidal dissipation in the shell without a direct requirement of a priori constraint on the shell’s rheology.
All studies N/A(62) N/A(28) 4 (11)
Park et al. (2024) 1 (87) 0.5 (40) 15 (20)
Park et al. (2024) 0.002 82 (134) 36 (61) 16 (21)
Park et al. (2024) 0.0002 116 (150) 51 (68) 22 (24)
Figure 7: Exploration of observables constraints from $J_{3}$, moment of inertia, libration amplitude, and rate of heat loss based on Park et al. (2024). We presented PDFs of predictions for $|k_{2}|$, $|h_{2}|$, $|l_{2}|$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$. Both scenarios of tidal dissipation are considered here. The plots of the PDFs for the $k_{2}$ amplitude and phase lag are notably different than those of the $|h_{2}|$ and $|l_{2}|$ amplitudes and phases. Plotting and unit conventions are the same as in Figure 4.

### v5

@@ 5 Road-map to Constrain the interior structure and tidal heating mode
Despite numerous studies based on Cassini observations, several interior structure properties of Enceladus are poorly constrained. This uncertainty is due in part to the limitation of the available observations and in part to differences in the interpretation of the data for the interior structure and the underlying modeling assumptions (e.g., Howett et al., 2011; Spencer et al., 2013; Thomas et al., 2016; Park et al., 2024; Schenk & McKinnon, 2024). Here, we explore the possible interior parameters that are consistent with the currently available observations. Based on these constraints, we predict ranges of the currently unavailable Love numbers, and investigate the possibility of discriminating between the two main tidal heating scenarios from future measurements.
@@
Currently available geophysical observations used to characterize Enceladus’s interior include shape (e.g., Park et al., 2024; Schenk & McKinnon, 2024), gravity up to degree-3 (Iess et al., 2014; Park et al., 2024), physical libration amplitude (Thomas et al., 2016; Nadezhdina et al., 2016; Park et al., 2024), and heat flux (Howett et al., 2011; Spencer et al., 2013) as summarized in Table 1. We note that the rates of heat loss considered here, except by Howett et al. (2011), are not based on direct measurements, but are estimated from the shell thickness and assumptions on its thermal conductivity (see discussions in Section 6). To avoid any self-consistency issues, we ensure that the obtained a posteriori ice shell thicknesses in our MCMC simulations are within the ranges used to determine the adopted rates of heat loss. We further discuss the assumptions on the rate of heat loss in Enceladus in Section 6.
@@
Figure 4: Exploration of the potentially observable geodetic quantities that are consistent with estimates of mass, heat flux (Howett et al., 2011; Hemingway & Mittal, 2019; Park et al., 2024), and libration amplitude spanned by estimates in Thomas et al. (2016), Nadezhdina et al. (2016) and Park et al. (2024). Other observables in this plot are free to vary. In the 2D Probability Density Functions (PDFs), black (yellow) color indicates low (high) probability. The observed multiple peaks in the PDFs of $|k_{2}|$, $|h_{2}|$, $|l_{2}|$, and phase lags reflect the different regimes of dissipation consistent with the observations (see Figure 3). Phase lags of $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, and $\epsilon_{l_{2}}$ are in degrees, amplitude of the libration is in meters, and $J_{3}$ is divided by 106.
5.1 Plausible ranges of observables and interior parameters
@@
The patterns of the posterior probability density functions (PDFs) of the amplitude and phase lag of $k_{2}$ largely differ from the PDFs of the amplitudes and phases of $h_{2}$ and $l_{2}$ (Figure 4), consistent with Section 4. PDFs of $|k_{2}|$ and $\epsilon_{k_{2}}$ have several peaks associated with different tidal dissipation mechanisms that can occur either in the shell or in the core. Each of these peaks correspond to the ranges of viscosities that can explain the estimated total heat shown in Figure 2. A different distribution pattern and peaks associated with different regimes of tidal activity are obtained for $h_{2}$ and $l_{2}$, and their phase lags. This different pattern arises from the fact that $|h_{2}|$ and $|l_{2}|$ are largely insensitive to tidal activity in the core, and thus do not reveal any peaks associated with the dissipation in the shell. Moreover, the posterior PDFs (Figure 4) show that while the $|h_{2}|$ and $|l_{2}|$ and their phase lags are strongly correlated with each other (dominated by the response of the shell). Their correlation with the amplitude and phase lag of $k_{2}$ depends on where the tidal dissipation dominantly occurs.
@@
We also obtain the ranges of the interior model parameters and the correlation between them associated with the observables chosen here (Appendix E). Previously, we ruled out mean crustal thicknesses around 5–10 km for Enceladus, as they would imply amplitudes of libration that are too large (see Section 3). Ice shell thicknesses $\approx$ 1–5 km would be compatible with the estimates of tidal heating and libration only if the mean effective viscosity of the shell is unrealistically low. Therefore, we conclude that the mean thickness of Enceladus’s shell is larger than 10 km. For completeness, we include very low effective viscosities for the shell and the core, which imply deformation in the viscous-dominated regimes. As a result, the PDFs of the viscosities of the shell and core indicate multiple peaks representative of tidal dissipation dominated by both viscous deformation as well as the shear (anelastic) relaxation (See also Figure 2). In the remaining analysis below, we consider viscosities of the ice shell and the rocky core that are close to or greater than their nominal relaxation viscosities ($\eta_{\rm shell}>10^{12}$ Pa.s and $\eta_{\rm core}>10^{11}$ Pa.s), i.e., the ice shell and the rocky core deform in the anelasticity-dominated deformation regime, as opposed to the viscous-dominated deformation regime.
@@
Next, we focus on the analysis of discriminating between two sets of geodetic observables that are associated with each of the two scenarios (Figure 5). In this figure, and in the remainder of this paper, we distinguish between core-dominated and shell-dominated tidal dissipation scenarios by associating them with the models where at least 90% of the total heat is produced in the shell or in the core. We compute the tidal heating in the shell using equation (9). The PDFs of $|k_{2}|$ and $\epsilon_{k_{2}}$ for the two tidal dissipation scenarios largely overlap, whereas the PDFs of $|h_{2}|$, $\epsilon_{h_{2}}$, $|l_{2}|$, and $\epsilon_{l_{2}}$ in the case of shell-dominated tidal dissipation differ significantly from those in the case of core-dominated tidal dissipation (Figure 5). In an inverse problem context, the degeneracy caused by the two different tidal dissipation scenarios reflected in the tidal Love numbers, as discussed earlier, leads to the non-uniqueness of the interpretation of $k_{2}$ and $\epsilon_{k_{2}}$.
@@
By differentiating between the two tidal dissipation scenarios, we obtain a tight constraint on the viscosity of the shell in the cases when the tidal dissipation is dominated in the shell. Similarly, we constrain the viscosity of the core in the case when we assume that tidal dissipation occurs predominantly in the core (Figure 6). Based on the set of observables used in Figures 4 and 5, we find that the thickness of the ice shell, mostly constrained by the libration amplitude, is in the range of 15–40 km. Because we have conservatively chosen broad ranges for observables in this analysis, we do not obtained tight constraints on most of the other model parameters in this analysis. The PDFs and correlations between the interior structure parameters are presented in Appendix E.
@@
Figure 5: Exploration of the potentially observable quantities consistent with estimates of mass, heat flux and libration amplitude for tidal dissipation scenarios in the shell and in the core. The heat flux is assumed as $\rm 15~GW<\dot{E}<40~GW$, covering the estimated ranges in Hemingway et al. (2018); Park et al. (2024), and the direct measurements (Howett et al., 2011). Libration is assumed to be between $\rm 0.082^{\circ}<\gamma<0.169^{\circ}$, covering the estimates in Thomas et al. (2016), Nadezhdina et al. (2016), and Park et al. (2024). Tidal heating produced in the shell in either scenario is also shown. In the 2D PDFs, the pink-gray and black-yellow themes refer to dissipation in the shell and and in the core, respectively. All variables are shown in logarithmic axes. Unit conventions are the same as in Figure 4.
@@
We compare the PDFs of $h_{2}$ and $l_{2}$ and their phase lags ($\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$) associated with each of the two heat generation scenarios to compute the minimum required precision of measurements to discriminate between the two dissipation scenarios (Figure 5). We compute the minimum required measurement precision of $|h_{2}|$ and $|l_{2}|$ by calculating the separation between the two intervals determined by the 2-$\sigma$ confidence regions associated with the two PDFs. Therefore, if we measure $h_{2}$ with a precision equal to the separation between the 2-$\sigma$ confidence intervals of the two PDFs, we obtain a confidence $\sim$95% to identify the layer that dominates the dissipation budget. Alternatively, if we measure $h_{2}$ with a precision equal to the separation between the 1-$\sigma$ confidence intervals, we obtain a confidence of $\sim$68% in identifying the layer dominating dissipation (not shown here). We present more detailed comparison of $|h_{2}|$ and $|l_{2}|$ and their phase lags in Appendix E.
@@
The PDFs of $|h_{2}|$ and $|l_{2}|$ associated with the two tidal dissipation scenarios using this set of observations demonstrate large overlaps (Figure 5), challenging the ability to distinguish between the dissipation in the shell or in the core. This large overlap is mainly due to the large uncertainties in the observables considered in this case. This analysis implies that measurement of the maximum vertical tidal deformations with a maximum uncertainty of 62 cm can distinguish between the dominant region of tidal activity within the 1-$\sigma$ confidence level. Measurement of radial deformations with a maximum uncertainty of 56 cm can indicates the “more likely” region of heat generation, but it cannot definitively discriminate between the two scenarios. Alternatively, measurement of the tangential displacements within an uncertainty of 28 cm is required for 1-$\sigma$ confidence. The phase lags associated with the radial deformations ($\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$) indicate a better separation between the PDFs associated with the two regimes of tidal activity compared to the deformations ($|h_{2}|$, $|l_{2}|$). We find that measuring the phase lag angles of the tidal deformations (either $\epsilon_{k_{2}}$ or $\epsilon_{l_{2}}$) with an accuracy of $1$ is required for 1-$\sigma$ confidence, and 4 is required for 2-$\sigma$ confidence in determining the dominant region of tidal dissipation (see Appendix E for more details).
5.2 Computing measurement requirements based on Park et al. (2024)
@@
Next, we focus on a narrower range of observational constraints by exclusively using the analysis by Park et al. (2024) for $\gamma$, $J_{3}$, moment of inertia (inferred from $C_{20}$ and $C_{22}$), total rate of heat loss (inferred from conductive shell thickness), and total density (see Table 1). We acknowledge that estimation of the moment of inertia in Park et al. (2024) is not completely independent of $\gamma$ and $J_{3}$. However, the estimated value of moment of inertia has been presented in a conservatively broad range and the effect of using it independently is negligible. We use 3-$\sigma$ uncertainties on all of observables. We present the correlation between these quantities incorporating both tidal dissipation scenarios in Figures 7. Applying the narrower observational constraints based on Park et al. (2024) shows that $|k_{2}|$ and $\epsilon_{k_{2}}$ indicate a normal distribution with a single peak that encompasses both scenarios of tidal dissipation, thus making them indistinguishable based only on measurement of $|k_{2}|$ and $\epsilon_{k_{2}}$. By contrast, $|h_{2}|$, $|l_{2}|$, $\epsilon_{h_{2}}$, and $\epsilon_{l_{2}}$ show two peaks associated with the two tidal dissipation scenarios. By separating the two tidal dissipation regimes (see figures in Appendix E), we also quantify the preferred values of the observables with the highest likelihood in the case of each of the two heat generation scenarios (Table 3). We note that satisfying all constraints provided in Park et al. (2024) is possible when using 3-$\sigma$ uncertainties on the observables. Specifically, we find that the mean value of $J_{3}$ can be matched with the libration amplitude when considering 3-$\sigma$ uncertainty in the measured $J_{3}$.
Our results can be compared to those by Genova et al. (2024) and Ermakov et al. (2021), where only tidal heating in the shell is considered. Also, Genova et al. (2024) use only the gravity measurements without exploiting the libration measurement as a constraint, which is expected to somewhat alter the results. The tighter constraints obtained based on Park et al. (2024) relax the precision requirements on the measurements of $h_{2}$ and $l_{2}$ required to discriminate between the two heating scenarios. Using the observational constraints of Park et al. (2024), we show the predicted amplitudes and phase lags of the Love numbers in Figure 8 panels (a)–(f). Using these narrower ranges of observational constraints, the PDFs of the tidal response associated with the two tidal dissipation scenarios are further separated compared to Section 5.1, where broader ranges of values for the observables were used (Figure 5). In this case, measuring the maximum vertical deformations with a precision of 87 cm is sufficient to differentiate between the two hypotheses of heat generation with a 1-$\sigma$ confidence. For a confidence level of 2-$\sigma$, measuring the maximum radial displacement of 1 cm is required. Alternatively, measuring the phase lag $\epsilon_{h_{2}}$ by 20 and 15 is sufficient to discriminate between the two tidal dissipation scenarios with 1-$\sigma$ and 2-$\sigma$ confidences, respectively. In addition, our analysis indicates that measurements of the maximum tangential displacements within 40 cm and 0.5 cm are required to differentiate between the two tidal dissipation scenarios providing 1-$\sigma$ and 2-$\sigma$ level confidences, respectively. The measurement precisions required for $\epsilon_{l_{2}}$ for the confidence levels of 1-$\sigma$ and 2-$\sigma$ are the same as those for $\epsilon_{h_{2}}$.
In case future measurements provide constraints on $k_{2}$, the measurement requirement for $h_{2}$ and/or $l_{2}$ could be relaxed. We present PDFs of displacement tidal Love numbers associated with the core- and shell-dominated tidal dissipation assuming constraints on the amplitude and phase lag of $k_{2}$ (Figure 8 panels (g)–(j)). If the real and imaginary parts of $k_{2}$ are measured within 0.002, the accuracies required to measure the maximum radial deformations to obtain 1-$\sigma$ and 2-$\sigma$ confidence levels relax to 134 cm and 82 cm, respectively. In that case, the upper bounds of the uncertainty required for the horizontal deformations are 61 cm and 36 cm, respectively. In this case, the accuracies required for the phase lags of $h_{2}$ and $l_{2}$ are equal to 21 and 16 for confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively.
In case more precise measurements of $k_{2}$ (both amplitude and phase) are achieved, i.e. within an uncertainty of 0.0002, the required measurement of radial and horizontal deformations will be further relaxed (not shown here). In that case, the requirement for the maximum radial displacement will be 150 cm and 116 cm for the 1-$\sigma$ and 2-$\sigma$ confidence levels, respectively. These values are 68 cm and 51 cm for the tangential component of the deformation, respectively. The precisions required for the phase lags of $h_{2}$ and $l_{2}$ are equal to 24 and $2$ to obtain confidence levels of 1-$\sigma$ and 2-$\sigma$, respectively. Due to the large overlap between the PDFs of $k_{2}$ between the two scenarios, we cannot differentiate between them even with high accuracy measurement of $k_{2}$ alone. We summarize the precision required for measuring the radial and horizontal displacements based on different scenarios discussed here in Table 4. Note that the measurements of the $h_{2}$ and $l_{2}$ as well as the $\epsilon_{h_{2}}$ and $\epsilon_{l_{2}}$ are correlated with each other and, therefore, measuring either of these quantities is sufficient for discriminating between tidal dissipation scenarios. We perform our analysis based on both the derivation in this paper (equation 11) and that in Beuthe (2019). We find that the difference in the results is less than 5%. This close approximation indicates the applicability of equation 11 in capturing the tidal dissipation in the shell without requiring a priori constraint on the shell’s rheology.
@@
All studies N/A(62) N/A(28) 4 (11)
Park et al. (2024) 1 (87) 0.5 (40) 15 (20)
Park et al. (2024) 0.002 82 (134) 36 (61) 16 (21)
Park et al. (2024) 0.0002 116 (150) 51 (68) 22 (24)
Figure 7: Exploration of observables constraints from $J_{3}$, moment of inertia, libration amplitude, and rate of heat loss based on Park et al. (2024). We present PDFs of predictions for $|k_{2}|$, $|h_{2}|$, $|l_{2}|$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$, $\epsilon_{l_{2}}$. Both scenarios of tidal dissipation are considered here. The plots of the PDFs for the $k_{2}$ amplitude and phase lag are notably different than those of the $|h_{2}|$ and $|l_{2}|$ amplitudes and phases. Plotting and unit conventions are the same as in Figure 4.
[…]
@@
Table 5: Ranges of the model parameters from two MCMC analysis satisfying the currently available measurements based on the analysis of observables in Park et al. (2024).
Using the ranges of $\gamma$, $J_{3}$, MoI, and heat inferred by Park et al. (2024), we also obtain considerably improved constraints on the interior parameters compared to the case where we use broader uncertainties on the observations (See Figures 18 and 21a in Appendix E). However, several model parameters remain weakly constrained by the currently available observations. The ranges of the obtained plausible interior parameters are listed in Table 5. We present PDFs of all the model parameters associated with each of the cases in Appendix E (Figures 21b and 21c).
To obtain further insights into the correlation between the observables and model parameters, we investigate the correlation between total heat and the observables and the interior parameters to which they are most sensitive (Figures 9a and 9b). We show the ranges of the effective viscosities of the shell and the core, as well as their correlation to the shell rigidity and $h_{2}$ and $k_{2}$ in Figure 9a. The total tidal heating is largely correlated to the viscosity of the region where the tidal dissipation is dominated. The rigidity of the shell influences the amplitude of both $k_{2}$ and $h_{2}$. We also explore the correlation between the physical properties of the shell (thickness, density, and rigidity) and the observables that are most sensitive to them, i.e., $\gamma$, $J_{3}$, and $k_{2}$ (Figure 9b). Shell thickness and density influence both $\gamma$ and $J_{3}$, although the effect of the density on $J_{3}$ is stronger than on $\gamma$. $D_{shell}$ also strongly correlates with $k_{2}$. The effect of the shell rigidity on the libration amplitude is small compared to the thickness and density which determine the shell’s moment of inertia. This sensitivity suggests that the shell’s response to the libration forces is mainly determined by inertial forces, consistent with the discussions in Section 3 and Appendix C and the results obtained by Hemingway & Nimmo (2024). The analysis shows that although libration is currently an effective constraint on the shell thickness, for obtaining very high accuracy on the shell and ocean thickness, precise measurements of the gravity field alongside the shape are more impactful. The lower sensitivity of libration is largely due to its dependence on the moment of inertia and flexibility of the shell which include effects from thickness, density, and rigidity of the shell.
5.3 Constraining the interior of Enceladus with future observations